OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=6, b=2, c=?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (mathstudent55):

Pythagoras theorem: c^2 = a^2 + b^2 c = sqrt(a^2 + b^2)

OpenStudy (ivettef365):

Same use pythagorean: C^2 = A^2 + b^2 c^2 = 6^2 + 2^2

OpenStudy (mathstudent55):

Just substitute your numbers into a and b, and solve for c.

OpenStudy (ivettef365):

c^2 = 36 + 4 c^2 = 40

OpenStudy (ivettef365):

therefore it s 2 square root of10

OpenStudy (anonymous):

my answer: 2 sq root 10 or ~ 6.325

OpenStudy (anonymous):

its 2 radical 10

OpenStudy (anonymous):

my answer wa right?!!!!!!!!

OpenStudy (anonymous):

Yup